A biochemist wants to make 500ml of 0.050M HCl by diluting a 6.0M HCl solution. How much of that solution should be used?


1
Expert's answer
2021-01-28T05:37:14-0500

C1V1=C2V2"\implies" V2= C1V1/V2= 0.500*0.050/6.0= 4.2 ml